OpenStudy (anonymous):

3x - 4y = 6 6x - 8y = 10 The system shown has _____ solution(s). A-no B-one C-infinite

OpenStudy (blurbendy):

i would use the elimnation method first

OpenStudy (anonymous):

How do I do that?

OpenStudy (blurbendy):

3x - 4y = 6 6x - 8y = 10 Let's eliminate x, so let's multiply the ENTIRE top equation by -2 -6x + 8y = -23 6x - 8y = 10 0 = -13 this is a false equation, so there are no solutions

OpenStudy (anonymous):

OKay hold on let me get paper;p

OpenStudy (blurbendy):

lol k

OpenStudy (anonymous):

So the top equation would be -6x-8y=-12?

OpenStudy (blurbendy):

it should be +8y -6x + 8y = -12 because we multiplied it by -2

OpenStudy (anonymous):

Ohh yeah my bad.

OpenStudy (blurbendy):

'sall good

OpenStudy (anonymous):

Okay(: So what's next?

OpenStudy (blurbendy):

when you add both equations, you get 0 = -13 which is a false statement because 0 doesnt equal -13. So there are no solutions.

OpenStudy (anonymous):

Ohh okay(: thanks.